### v2

@@ 2.3 Positioning of Our Work
The discussions above have been exclusively focused on APR, as our method falls under this category as in Table 1. In terms of alternative techniques related to camera localization, we have classified them into several principal categories. Unlike other approaches, our strategy is distinct in that it eliminates the need for pre-built 3D models, initial pose priors, or iterative refinement during the pose estimation of new images.
@@ 3 Method
Figure 2: Overview. The training stage consists of two modules: APRNet, which handles camera pose regression and image feature extraction, and NeRF-P, which is responsible for view synthesis and volumetric pose feature extraction. The testing stage, which uses a simplified version of APRNet for fast inference, is highlighted in the blue dotted box.
@@ 3.1 Camera Localization with PoseMap
NeRF-P includes two distinct branches: the rendering branch and the volumetric pose extraction branch for generating PoseMap, which is the key component of our design. The details will be described in Sec 3.2.
@@ 4.2 Evaluation on Datasets
Evaluation on 7-scenes dataset. First, we compare our methods with one-frame APR methods on 7-scenes dataset. Tab. 2 shows that our method achieves the best results on all indoor scenes. To be specific, our main pipeline, denoted as PMNet, achieves state-of-the-art performance. Moreover, PMNetud finetuned in the self-supervising stage with unlabelled data further enhances the pose regression accuracy. The translation and rotation errors are reduced by 14.28% (0.07$\rightarrow$0.06) and 12.67%(2.21$\rightarrow$1.93) on average, respectively. For the Heads and Pumpkin scenes, our results are comparable to those of DFNet, though not superior. Upon reviewing the dataset, we found that these scenes are relatively small, with limited visible range and camera pose variations. Consequently, the benefits of our PoseMap may not be particularly significant.
@@
Evaluation on Cambridge landmarks dataset. Then we evaluate our methods on the more challenging dataset, i.e., Cambridge Landmarks. As shown in Tab. 3, our PMNet leads to the dominant performance advantage over all scenes in single-frame APR methods. Compared to DFNetdm [7] which also uses unlabelled data for online training, our method gets performance gain by 20.51% (0.39$\rightarrow$0.31) and 17.71% (0.96$\rightarrow$0.79) for translations and rotations, respectively.
@@ 4.3 Ablation Studies
Design choices of PoseMap. Thanks to the neural volumetric pose feature, aka PoseMap, our method achieves more accurate estimation results than previous methods. The underlying strength of PoseMap lies in its ability to establish stable geometric correspondences between the image and the camera pose across diverse scenes. To demonstrate the impact of PoseMap, we break down the pose features into three configurations in Table 5(a). The baseline corresponds to the default settings of DFNet [7] without any pose features supervision. $L_{nerfmap}$ refers to using the feature embedding extracted from the last MLP layer of pret-rained NeRF-H by directly rendering a feature map (termed as NeRFMap) for APRNet supervision. The comparatively weak results indicate that the original NeRF features do not inherently encapsulate camera pose information. By incorporating a pose branch and a decoder, without modifying the original NeRF training protocol, PoseMap is able to learn the implicit features from the ground truth pose. With the assistance of PoseMap and the $L_{posemap}$ loss.
@@
Visualization of PoseMap. In Fig 5, we visualize APR image features and PoseMap features. The APR image features are extracted from the output before the fully-connected layer of VGG16 network. Compared with APR image features, the proposed PoseMap features capture implicit features of the camera pose by aggregating global attributes from samples of light rays, resulting in local features with clearer geometric information than those from 2D-CNN backbones, which is useful on camera localization tasks as proven structure-based methods.
@@ 4.4 Discussions
Robustness of training size. Since the effectiveness of camera localization correlates with the scale of the observed images [34, 26], we perform an ablation experiment on the quantity of training data. We train the whole pipeline with 100%, 50%, and 25% of the training set in Kings of Cambridge Landmarks dataset, the evaluations are shown in Fig 6. As the data size decreases, the original DFNet[7] exhibits a rapid decrease in accuracy, indicating its high sensitivity to the quantity of input data. By comparison, our results are more robust to the size of training data as well as better performance.

### v3

@@ Abstract
Keywords:
Absolute pose regression Pose feature NeRF
@@ 2.3 Positioning of Our Work
The discussions above have been exclusively focused on APR, as our method falls under this category as in Table 1. In terms of alternative techniques related to camera localization, we have classified them into several principal categories. We emphasize that APR methods can deduce camera poses from unseen images using a single inference pass of a network, obviating the need for 2D-2D/3D matching with pre-built models or pose priors.
@@ 3 Method
Figure 2: Overview of the camera localization pipeline with PoseMap. The training stage of our pipeline, including two main modules: APRNet for camera pose regression and extracting image features and NeRF-P for view synthesis and extracting pose features. The inference stage of our pipeline with a simple APRNet for fast inference is highlighted in blue dotted box.
@@ 3.1 Camera Localization with PoseMap
NeRF-P includes two distinct branches: the rendering branch and the volumetric pose extraction branch for generating PoseMap, which is the key component of our design. Unlike CROSSFIRE [25] and NeFeS [6] that extend NeRF to synthesize CNN image features, we focuses on establishing a direct relationship between implicit 3D information and camera poses. The details will be described in Sec. 3.2.
@@ 4.2 Evaluation on Datasets
Evaluation on 7-scenes dataset. First, we compare our methods with one-frame APR methods on 7-scenes dataset. Table 2 shows that our method achieves the best results on all indoor scenes. To be specific, our main pipeline, denoted as PMNet, achieves state-of-the-art performance. Moreover, PMNetud finetuned in the self-supervising stage with unlabelled data further enhances the pose regression accuracy. The translation and rotation errors are reduced by 14.28% (0.07$\rightarrow$0.06) and 12.67%(2.21$\rightarrow$1.93) on average, respectively. For the Heads and Pumpkin scenes, our results are comparable to those of DFNet, though not superior. Upon reviewing the dataset, we found that these scenes are relatively small, with limited visible range and camera pose variations. Consequently, the benefits of our PoseMap may not be particularly significant.
@@
Evaluation on Cambridge landmarks dataset. Then we evaluate our methods on the more challenging dataset, i.e., Cambridge Landmarks. As shown in Table 3, our PMNet leads to the dominant performance advantage over all scenes in single-frame APR methods. Compared to DFNetdm [7] which also uses unlabelled data for online training, our method gets performance gain by 20.51% (0.39$\rightarrow$0.31) and 17.71% (0.96$\rightarrow$0.79) for translations and rotations, respectively.
@@ 4.3 Ablation Studies
Design choices of PoseMap. Thanks to the neural volumetric pose feature, aka PoseMap, our method achieves more accurate estimation results than previous methods. The underlying strength of PoseMap lies in its ability to establish stable geometric correspondences between the image and the camera pose across diverse scenes. To demonstrate the impact of PoseMap, we break down the pose features into three configurations in Table 5(a). The baseline corresponds to the default settings of DFNet [7] without any pose features supervision. $L_{nerfmap}$ refers to using the feature embedding extracted from the last MLP layer of pre-trained NeRF-H by directly rendering a feature map (termed as NeRFMap) for APRNet supervision. The comparatively weak results indicate that the original NeRF features do not inherently encapsulate camera pose information. By incorporating a pose branch and a decoder, without modifying the original NeRF training protocol, PoseMap is able to learn the implicit features from the ground truth pose. With the assistance of PoseMap and the $L_{posemap}$ loss.
@@
Visualization of PoseMap. In Fig. 5, we visualize APR image features and PoseMap features. The APR image features are extracted from the output before the fully-connected layer of VGG16 network. Compared with APR image features, the proposed PoseMap features capture implicit features of the camera pose by aggregating global attributes from samples of light rays, resulting in local features with clearer geometric information than those from 2D-CNN backbones, which is useful on camera localization tasks as proven structure-based methods.
@@ 4.4 Discussions
Robustness of training size. Since the effectiveness of camera localization correlates with the scale of the observed images [34, 26], we perform an ablation experiment on the quantity of training data. We train the whole pipeline with 100%, 50%, and 25% of the training set in Kings of Cambridge Landmarks dataset, the evaluations are shown in Fig. 6. As the data size decreases, the original DFNet[7] exhibits a rapid decrease in accuracy, indicating its high sensitivity to the quantity of input data. By comparison, our results are more robust to the size of training data as well as better performance.
@@ 5 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (62072422).
6 Supplementary
In the supplementary materials, we provide further insights of PoseMap designs in Sec. 6.1, implementation details in Sec. 6.2, extended discussions in Sec. 6.3, and visualizations in Sec. 6.4.
6.1 Technical Details of PoseMap
Our motivation is to improve the performance of camera localization with a novel feature map called PoseMap. Considering that NeRF itself takes the camera pose as input and conducts a volumetric rendering process. We surmise that it can help to identify an appropriate implicit feature representation for the camera pose.
We build our PoseMap based on the NeRF-W [22] and NeRF-H [7]. These two methods are designed for synthesizing novel views of complex scenes using only unstructured collections of in-the-wild photographs with different foreground occlusions and exposure levels. NeRF-W models the scene as a combination of both shared and image-specific elements, which allows for the unsupervised separation of scene’s content into static and transient components. In our exploration of pose features, we recognize that the camera pose information is more linked with the static background. Thus, we design our pose branch to depend on these static elements. Within this branch, the pose embedding is then thoroughly distilled from the ground truth camera pose followed by a series of MLPs. To be specific, $\mathcal{F}_{image}$ comes from three CNN layers at $(\frac{H}{2},\frac{W}{2},128)$. $\mathcal{F}_{posemap}$ is rendered from the pose branch of NeRF-P at $(\frac{H}{16},\frac{W}{16},256)$. Both are pixel-wise features.
In order to explore the implicit information of PoseMap, We compare it with NeRFMap which is directly extracted from the original NeRF. As rendered and visualized in Fig. 7, the salient region of PoseMap is more related to the geometrical structures such as lines and boundaries, which provides the obvious low-level evidence for camera poses estimation.
Figure 7: Visualization comparison between NeRFMap and PoseMap. From left to right, we show the RGB image (left), the NeRFMap (the middle column), and our PoseMap (right).
6.2 Implementation Details
6.2.1 Datasets
The numbers of frames of each scene in 7-scenes [39] and Cambridge Landmarks [17] dataset are listed in Table 6. For a fair comparison with other methods, we follow the sampling strategy in DFNet to use subsampled training data with a spacing window $d=2$ in Cambridge Landmarks, $d=5$ in 7-scenes for scenes containing $\leq$ 2000 frames and $d=10$ otherwise. The ground-truth pose annotated in 7-scenes and Cambridge Landmarks are obtained from RGB-D SLAM method named dSLAM [30] and COLMAP [35],
Table 6: The metadata of 7-scenes [39] and Cambridge Landmarks [17]
Scene Chess Fire Heads Office Pumpkin Kitchen Stairs
train 4000 2000 1000 6000 4000 7000 2000
test 2000 2000 1000 4000 2000 5000 1000
Scene Kings Hospital Shop Church
train 1220 895 231 1487
test 343 182 103 530
Table 7: The training time and memory cost of different stages.
Stages NeRF-P PMNet (APR / NeRF-P) PMNetud (APR / NeRF-P)
Duration(hrs) 13.0 18.1 (7.5 / 10.6) 27.9 (11.1 / 16.8)
Memory(GB) $\sim$5.0 $\sim$12.8 $\sim$18.4
6.2.2 Training time and memory cost
We report the training time and memory cost of Kings from Cambridge Landmarks. The overall training takes about 2.4 days on a single NVIDIA RTX 4090 GPU, including NeRF, PMNet, PMNetud. To be specific, APR training time makes up around $31.5\%$(18.6hrs), the training and rendering of NeRF account for nearly $68.5\%$(40.4hrs). Notably, NeRF-based methods (DFNet, LENS, etc.) have extended training duration due to extensive computational burden of NeRF itself. We plan to opt for more efficient rendering schemes in the future. However, inference time is not affected. Besides, excluding NeRF’s rendering time, the APR training time for PMNet (7.5 hrs) slightly exceeds that of DFNet (6.0 hrs) due to the computational demands of PoseMap. We firmly believe the notable improvement in performance, as in quantitative and qualitative results, outweighs the slight increase in computational resources.
6.2.3 Inference
For the inference time, our method runs at 1.82ms (about 550Hz) for a $240\times 427$ image and 6.27ms (about 160Hz) for a $480\times 854$ image. As for training duration, since we have incorporated NeRF into our online training pipeline, the time required for training is largely dominated by the volumetric rendering process. All our experiments are conducted on an NVIDIA GeForce RTX 4090 GPU with PyTorch.
6.3 Discussions
6.3.1 Comparing with refinement methods
Unlike APR-based methods that directly estimate the camera pose, the iterative refinement-based methods require prior estimation of the camera pose and continuously refine it based on the consistency of pixel features. Generally, these methods can achieve higher accuracy in pose estimation. In the experiments, as shown in Table 3 of the paper, our approach, PMNetud, achieves comparable accuracy of pose estimation to refinement-based methods. This suggests that, due to the design of our PoseMap and the domain consistency in implicit pose features, our algorithm can effectively enhance the precision of APR-based methodologies. In certain scenarios, it even outperforms refinement-based methods, demonstrating greater practical utility.
6.3.2 Comparing with SfM///SLAM methods
As illustrated in Sec. 6.2.1, most datasets of camera localization are annotated by SfM or SLAM. SfM aims to recover 3D geometry and camera parameters from image sets, which is done classically by computing local image features, finding matches across images and then estimating and verifying epipolar geometry using bundle adjustment. SLAM is used for localizing and mapping the surroundings from a video stream by matching feature descriptors in adjacent frames. As noted in Table 1 of the paper, the main differences between SfM-based methods and APR-based methods are two-fold. First, SfM-based methods are the stage of in-sample building with a set of in-sample images with unknown poses, which is built to obtain the poses of in-sample images. However, APR-based methods are in the stage of out-of-sample testing, aiming to use in-sample images to train a practical network for out-of-sample camera pose regression. Second, as the ground truth is obtained with COLMAP (over all frames of each scene) in the Cambridge Landmarks dataset, we believe the accuracy of all APR related methods would be limited by this upper bound. Therefore, we claim that it’s unfair to evaluate the performance comparisons with these methods.
6.4 Visualization
Here, we present more visualization samples of PoseMap in Fig. 8. As we can see, APR feature map is a CNN-based feature map pre-trained on ImageNet datasets. The salient region is more related to the semantics and appearance. However, our PoseMap is an MLP-based feature map created through volumetric rendering, which captures geometric details more effectively for improved pose estimation. Additionally, the supervision of APR feature map faces the challenge of a domain gap between real images and rendered images, an issue that doesn’t apply to the PoseMap.
Figure 8: Visualization of more localization results. We show (a) the input real image, (b) the rendered image using the pose estimated by PMNetud, (c) the APR feature map and (d) our PoseMap. Note that, The dimensionality reduction via PCA is used to visualize the feature map.